OpenStudy (anonymous):

3 square root -540x^15y^20 raise to the power or take the root. Assume that all variables represent positive numbers?????

OpenStudy (saifoo.khan):

\[\sqrt[3]{-540x^{15}y^{12}}\]This right?

OpenStudy (cwrw238):

i think its y^20 saifoo

OpenStudy (saifoo.khan):

Ops! Right. My bad. typo.

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{-540x^{15} y^{20}} \to -3x^5 y^6 \sqrt[3]{20y^2}\]

OpenStudy (anonymous):

Thank you.